calculate the molar mass of (NH4)2SO4
1
Expert's answer
2020-02-03T07:45:54-0500

М((NH4)2SO4) = 2*Ar(N) + 4*2*Ar(H) + Ar(S) + 4*Ar(O) = 2*14 + 8*1 + 32 + 4*16 = 132(g/mol)